A line with slope -2 passes through the point (1,4). Find the area of the triangle formed by this line and the x and y axes.
Answer by lwsshak3(11628)

A line with slope -2 passes through the point (1,4). Find the area of the triangle formed by this line and the x and y axes.
***
Equation of a line: y=mx+b, m=slope, b=y-intercept
y=-2x+b
solving for b
4=-2+b
b=6
equation of given line: y=-2x+6
..
set y=0
-2x+6=0
x=3 (x-intercept=base of triangle)
y=6 (y-intercept=height of triangle)
area of triangle=1/2*base*height=1/2*3*6=9 sq units